12. The only rocket that has ever lifted humans beyond low Earth orbit was the Saturn V, which had three stages of propulsion. Fully fueled, the rocket plus its cargo had a mass of 2.97x10 kg. The first stage consumed fueltoxidizer at a rate of 1.5x10 kgs, and had an exhaust speed of 2.6x10*m/s. (a) Calculate the thrust that was produced by the first stage. (b) Determine the acceleration of the rocket just after lift-off from the launch pad. 3
Answer #1

Given,

M = 2.97 x 10^6 kg ; dm/dt = 1.5 x 10^4 kg/s ; v = 2.6 x 10^3 m/s

a)We know that

T = dm/dt x v

T = 1.5 x 10^4 x 2.6 x 10^3 = 3.9 x 10^7 kg- m/s^2

Hence, T = 3.9 x 10^7 kg-m/s^2 or N

b)T = M a=> a = T/M

a = 3.9 x 10^7/2.97 x 10^6 = 13.13 m/s^2

Hence, a = 13.13 m/s^2